### v1

# Wilcoxon-Mann-Whitney Test of No Group Discrimination

Abstract.
The traditional WMW null hypothesis $\mathrm{H_{0}}\textrm{:}\,\mathrm{F=G}$ is erroneously too broad. WMW actually tests narrower $\mathrm{H_{0}}\textrm{:}\,\textrm{AUC}=0.5$. Asymptotic distribution of the standardized $U$ statistic (i.e., the empirical AUC) under the correct $\mathrm{H_{0}}$ is derived along with finite sample bias corrections. The traditional alternative hypothesis of stochastic dominance is too narrow. WMW is consistent against $\mathrm{H_{1}}\textrm{:}\,\textrm{AUC}\neq 0.5$, as established by Van Dantzig in 1951.
1. Introduction
The Wilcoxon-Mann-Whitney (WMW) test statistic [wilcoxon1945individual, mann1947test] is the $U$-statistic $U=\sum_{i}\sum_{j}1\{X_{i}>Y_{j}\}$. Its standardized version $U/(n_{1}n_{2})$ [birnbaum1955] is the empirical AUC ($\mathrm{eAUC}=\hat{P}(X>Y)$), cf. [pepe2003statistical], connecting the familiar rank-based framework to discrimination analysis.
WMW is traditionally stated to test $\mathrm{H_{0}}\textrm{:}\,\mathrm{F=G}$ against $\mathrm{H_{1}}\textrm{:}\,\mathrm{F\neq G}$ or alternatives of stochastic dominance. This formulation appears in most statistical textbooks [lehmann1975nonparametrics] and software implementations [R]. We argue this formulation is incorrect.
2. Theoretical Evidence Against Traditional $\mathrm{H_{0}}$
Consider two independent samples from $N(0,\sigma_{1}^{2})$ and $N(0,\sigma_{2}^{2})$ with $\sigma_{1}\neq\sigma_{2}$. Here $\mathrm{F\neq G}$ due to different variances, so the traditional null hypothesis $\mathrm{H_{0}}\textrm{:}\,\mathrm{F=G}$ is false and $\mathrm{H_{1}}\textrm{:}\,\mathrm{F\neq G}$ holds.
However, Monte Carlo simulation demonstrates that the test statistic eAUC concentrates asymptotically on 0.5 - the value expected under a true null hypothesis. The simulation (10,000 replications, $\sigma_{1}=0.1$, $\sigma_{2}=3$, $n_{1}=n_{2}=n=1000$) yields empirical mean $0.5000069$ and $\mathrm{SD}\,\,0.01554$. This agrees closely with the theoretical asymptotic distribution $N(1/2,1/(4n))$ derived under $\mathrm{H_{0}}\textrm{:}\,\textrm{AUC}=0.5$, which predicts $\mathrm{SD}\,\,=0.01581$ for $n=1000$.
The empirical AUC concentrates precisely on 0.5 despite the traditional $\mathrm{H_{0}}\textrm{:}\,\mathrm{F=G}$ being false. This creates a logical contradiction: if WMW tested distributional equality, the test statistic should not concentrate on its null value when distributions clearly differ.
Though the Gaussian example invalidates the traditional claim that WMW tests $\mathrm{H_{0}}\textrm{:}\,\mathrm{F=G}$, it does not specify the maximal class of alternatives that WMW can detect. This is provided by Van Dantzig [vandantzig1951] who proved that WMW is consistent against alternatives $\mathrm{H_{1}}\textrm{:}\,\textrm{AUC}\neq 0.5$, where $\mathrm{AUC}=P(X>Y)$ for independent observations from the two populations. For sufficiently small significance levels, the test is consistent against no other alternatives. This consistency characterization establishes that WMW tests $\mathrm{H_{0}}\textrm{:}\,\textrm{AUC}=0.5$, detecting systematic pairwise dominance rather than general distributional differences.
3. The Most General Alternative Hypothesis of WMW
Traditional alternative hypothesis of WMW test is $\mathrm{H_{1}}\textrm{:}\,\mathrm{F\neq G}$, cf. [lehmann1975nonparametrics]. Lehmann [lehmann1975nonparametrics] notes that not all pairs of distributions $\mathrm{F\neq G}$ are of equal relevance, and then defines stochastic dominance as the alternative hypothesis of interest. This restriction is unnecessarily narrow. Van Dantzig’s consistency result shows WMW detects any departure from $\mathrm{AUC}=0.5$. This includes cases where distributions cross multiple times, violating stochastic dominance assumptions, yet systematic pairwise advantage exists. The correct alternative $\mathrm{H_{1}:AUC}\neq 0.5$ subsumes stochastic dominance as a special case while capturing the full range of discrimination patterns WMW can detect.
4. Asymptotic Distribution of Empirical AUC Under the Null Hypothesis $\mathrm{H_{0}}\textrm{:}\,\textrm{AUC}=0.5$
4.1. General Framework
Consider the general setup with identically and independently distributed (i.i.d.) random variables taking values in $\mathbb{R}$: $X_{1},\ldots,X_{n_{1}}\overset{\text{i.i.d.}}{\sim}F$, $Y_{1},\ldots,Y_{n_{2}}\overset{\text{i.i.d.}}{\sim}G$. Let $n=n_{1}+n_{2}$, $\lambda_{n}=\frac{n_{1}}{n}\to\lambda\in(0,1)$. $A=\int F\,dG$ is $0.5$ by assumption. And $\hat{A}$ is an estimator of AA
$\hat{A}=(n_{1}n_{2})^{-1}\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}1\{X_{i}<Y_{j}\}.$
4.2. U-statistic Central Limit Theorem
From the general two-sample U-statistic CLT:
$\sqrt{n}(\hat{A}-A)\xrightarrow{d}\mathcal{N}(0,\sigma^{2}),\quad\sigma^{2}=\frac{\zeta_{1}^{2}}{\lambda}+\frac{\zeta_{2}^{2}}{1-\lambda},$
where
(1) $\zeta_{1}^{2}$ $=\text{Var}(G(X))=\int G^{2}\,dF-\left(\int G\,dF\right)^{2},$
$\zeta_{2}^{2}$ $=\text{Var}(F(Y))=\int F^{2}\,dG-\left(\int F\,dG\right)^{2}.$
Under the assumption, $\int G\,dF=0.5$ and $\int F\,dG=0.5$, so (1) simplifies to
$\zeta_{1}^{2}$ $=\int G^{2}\,dF-0.25,$
$\zeta_{2}^{2}$ $=\int F^{2}\,dG-0.25.$
4.3. Placement Variables
Define the placement variables
$U:=G(X)\sim H_{1},\quad V:=F(Y)\sim H_{2}.$
Since $F$ and $G$ are continuous, $U$ and $V$ are uniform on $[0,1]$ if and only if $F=G$. In general, they follow distributions $H_{1},H_{2}$ on $[0,1]$ satisfying
$E[U]=0.5,\quad E[V]=0.5.$
Then
$\zeta_{1}^{2}$ $=\text{Var}(U)=E[U^{2}]-0.25,$
$\zeta_{2}^{2}$ $=\text{Var}(V)=E[V^{2}]-0.25.$
Hence
$\sigma^{2}=\frac{\text{Var}(U)}{\lambda}+\frac{\text{Var}(V)}{1-\lambda}.$
$4.$ Asymptotic Distribution
Under the restriction $\int F\,dG=0.5$:
$\sqrt{n}(\hat{A}-0.5)$ $\xrightarrow{d}\mathcal{N}(0,\sigma^{2}),$
$\sigma^{2}$ $=\frac{\text{Var}(G(X))}{\lambda}+\frac{\text{Var}(F(Y))}{1-\lambda}.$
4.5. Consistent Variance Estimation
Compute the empirical placement values
$\hat{G}(X_{i})$ $=n_{2}^{-1}\sum_{j=1}^{n_{2}}1\{Y_{j}\leq X_{i}\},$
$\hat{F}(Y_{j})$ $=n_{1}^{-1}\sum_{i=1}^{n_{1}}1\{X_{i}\leq Y_{j}\}.$
Then
$\widehat{\text{Var}}(G(X))$ $=(n_{1}-1)^{-1}\sum_{i=1}^{n_{1}}(\hat{G}(X_{i})-0.5)^{2},$
$\widehat{\text{Var}}(F(Y))$ $=(n_{2}-1)^{-1}\sum_{j=1}^{n_{2}}(\hat{F}(Y_{j})-0.5)^{2},$
and
$\hat{\sigma}^{2}=\frac{\widehat{\text{Var}}(G(X))}{\lambda_{n}}+\frac{\widehat{\text{Var}}(F(Y))}{1-\lambda_{n}}$
is consistent for $\sigma^{2}$.
Therefore,
$\hat{A}\pm z_{\alpha/2}\cdot\frac{\hat{\sigma}}{\sqrt{n}}$
provides an asymptotically valid $(1-\alpha)$ confidence interval for $A=0.5$.
4.6. Finite Sample Bias Correction
Under the restriction $A=\int F\,dG=0.5$ but $F\neq G$, we derive bias correction using general U-statistic theory.
4.6.1. Bias of Naive Estimators
Let $\zeta_{1}^{2}=\text{Var}(G(X))$ and $\hat{\zeta}_{1}^{2}=(n_{1}-1)^{-1}\sum_{i=1}^{n_{1}}(\hat{G}(X_{i})-0.5)^{2}$.
[…]
$\hat{G}(x)=G(x)+\Delta_{2}(x),\quad\text{where }\Delta_{2}(x)=\frac{\sqrt{R_{2}(x)}}{n_{2}}\text{ and }R_{2}(x)\approx G(x)(1-G(x))\cdot n_{2}.$
A first-order stochastic expansion gives
$E[\hat{\zeta}_{1}^{2}]=\zeta_{1}^{2}-\frac{1}{n_{2}}E[G(X)(1-G(X))]+o(1/n_{2}).$
Therefore, the finite-sample bias is
$\text{Bias}_{1}=-\frac{1}{n_{2}}E[G(X)(1-G(X))].$
An analogous calculation for the second term yields
$\text{Bias}_{2}=-\frac{1}{n_{1}}E[F(Y)(1-F(Y))].$
4.6.2. Bias-Corrected Estimators
Estimate the expectations by the plug-in principle:
$\hat{\omega}_{1}$ $=\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\hat{G}(X_{i})(1-\hat{G}(X_{i})),$
$\hat{\omega}_{2}$ $=\frac{1}{n_{2}}\sum_{j=1}^{n_{2}}\hat{F}(Y_{j})(1-\hat{F}(Y_{j})).$
The bias-corrected estimators are
[…]
4.6.3. Final Variance Estimator
Insert (2) into the Welch-type combination:
$\hat{\sigma}^{2}_{\text{adj}}=\frac{\widehat{\text{Var}}^{*}(G(X))}{\lambda_{n}}+\frac{\widehat{\text{Var}}^{*}(F(Y))}{1-\lambda_{n}},$
with $\lambda_{n}=n_{1}/n$.
Use $\hat{\sigma}^{2}_{\text{adj}}$ together with Welch-Satterthwaite degrees of freedom for the combined variance estimator:
$\text{df}=\frac{(\hat{\sigma}^{2}_{\text{adj}})^{2}}{\frac{[\widehat{\text{Var}}^{*}(G(X))/\lambda_{n}]^{2}}{n_{1}-1}+\frac{[\widehat{\text{Var}}^{*}(F(Y))/(1-\lambda_{n})]^{2}}{n_{2}-1}}$
and the $t$-quantile $t_{\alpha/2,\text{df}}$ to obtain
$\hat{A}\pm t_{\alpha/2,\text{df}}\cdot\frac{\hat{\sigma}_{\text{adj}}}{\sqrt{n}}.$
4.7. Higher-Order Correction
For enhanced finite-sample performance, a data-based correction that reduces remaining bias:
(3) $\hat{\sigma}^{2}_{\text{final}}=\left(1-\frac{1}{n_{1}}-\frac{1}{n_{2}}\right)\cdot\hat{\sigma}^{2}_{\text{adj}}.$
Equation (3) provides $O(n^{-2})$ coverage accuracy for any continuous $F,G$ with $A=0.5$.
4.7.1. Summary
Whenever the true AUC is 0.5, the empirical AUC is asymptotically normal:
(1) $\sqrt{n}(\hat{A}-0.5)\xrightarrow{d}\mathcal{N}(0,\sigma^{2}),\quad\sigma^{2}=\frac{\text{Var}(G(X))}{\lambda}+\frac{\text{Var}(F(Y))}{1-\lambda},$
with consistently estimable variance via the bias-corrected estimators.
4.8. R implementation
Traditional asymptotic p-values are derived under $\mathrm{H_{0}}\textrm{:}\,\mathrm{F=G}$ and employ variance estimation appropriate for that hypothesis. However, as demonstrated above, WMW actually tests $\mathrm{H_{0}}\textrm{:}\,\textrm{AUC}=0.5$. This discrepancy leads to miscalibrated inference, particularly under heteroscedasticity.
Even in the location-shift special case where $\mathrm{H_{0}}\textrm{:}\,\mathrm{F=G}$ $\Leftrightarrow$ $\mathrm{H_{0}}\textrm{:}\,\textrm{AUC}=0.5$ the variance estimation under the $\mathrm{F=G}$ framework can differ from that under the correct $\mathrm{AUC=0.5}$ framework, leading to different p-values.
The correct asymptotic p-values and finite-sample bias corrections derived under $\mathrm{H_{0}}\textrm{:}\,\textrm{AUC}=0.5$ are implemented in the R package wmwAUC, available from [grendar2025repo].
5. Notes
For historical precision, Van Dantzig [vandantzig1951] formulated the alternative hypothesis in terminology that would today be stated as ’the Area above the Ordinal Dominance Graph $\neq 0.5$’. The Area above the Ordinal Dominance Graph [bamber1975area] corresponds to the AUC. For application of the dominance statistic in psychology, see [cliff1993dominance].
Notably, Van Dantzig [vandantzig1951] paired his AUC-based alternative hypothesis $\mathrm{H_{1}}\textrm{:}\,\textrm{AUC}\neq 0.5$ with the traditional null hypothesis $\mathrm{H_{0}}\textrm{:}\,\mathrm{F=G}$. This creates a logical inconsistency, as $\mathrm{H_{0}}\textrm{:}\,\mathrm{F=G}$ is not the proper complement to $\mathrm{H_{1}}\textrm{:}\,\textrm{AUC}\neq 0.5$. The correct pairing is $\mathrm{H_{0}}\textrm{:}\,\textrm{AUC}=0.5$ versus $\mathrm{H_{1}}\textrm{:}\,\textrm{AUC}\neq 0.5$.
The condition $P(X\leq Y)\geq 1/2$ is defined in [arcones2002nonparametric] as stochastic precedence.
Under location-shift assumption, the traditional $\mathrm{H_{0}}\textrm{:}\,\mathrm{F=G}$ and the correct $\mathrm{H_{0}}\textrm{:}\,\textrm{AUC}=0.5$ are mathematically equivalent. However, the asymptotic distributions used to derive p-values differ between the two frameworks, leading to different variance estimation and potentially different inference.
While this work focuses on continuous random variables, the core critique of traditional hypothesis formulations extends to discrete cases, though the positive characterization of what WMW tests becomes more complex due to tied observations. Similar issues arise in k-sample extensions (Kruskal-Wallis test), though the correct characterization requires investigation beyond the present scope.
Recent work has begun addressing limitations of traditional nonparametric frameworks. Conroy [conroy2012hypotheses] highlighted misinterpretations of WMW as testing median equality, emphasizing its role as measuring $P(X>Y)$. In [del2025invariant], the authors developed measures of disagreement when stochastic dominance assumptions fail. The present work provides a more fundamental resolution by establishing the correct null hypothesis and complete characterization of WMW sensitivity.
Acknowledgments
Valuable discussions with Claude AI Sonnet 4 (Anthropic, 2025) and Kimi AI (Moonshot AI, 2025), which contributed to refining ideas presented in this work, are gratefully acknowledged. After using these tools the author reviewed and edited the content as necessary and takes full responsibility for the content of the publication.