### v2

# Wilcoxon-Mann-Whitney Test of No Group Discrimination

Abstract.
The traditional WMW null hypothesis $\mathrm{H_{0}\colon F=G}$ is erroneously too broad. WMW actually tests narrower $\mathrm{H_{0}\colon AUC=0.5}$. Asymptotic distribution of the standardized $U$ statistic (i.e., the empirical AUC) under the correct $\mathrm{H_{0}}$ is derived along with finite sample bias corrections. The traditional alternative hypothesis of stochastic dominance is too narrow. WMW is consistent against $\mathrm{H_{1}\colon AUC\neq 0.5}$, as established by Van Dantzig in 1951.
1. Introduction
The Wilcoxon-Mann-Whitney (WMW) test statistic [wilcoxon1945individual, mann1947test] is the $U$-statistic $U=\sum_{i}\sum_{j}1\{X_{i}<Y_{j}\}$. Its standardized version $U/(n_{1}n_{2})$ [birnbaum1955] is the empirical AUC ($\mathrm{eAUC}=\hat{P}(X<Y)$), cf. [pepe2003statistical], connecting the familiar rank-based framework to discrimination analysis.
WMW is traditionally stated to test $\mathrm{H_{0}\colon F=G}$ against $\mathrm{H_{1}\colon F\neq G}$ or alternatives of stochastic dominance. This formulation appears in most statistical textbooks [lehmann1975nonparametrics] and software implementations [R]. We argue this formulation is incorrect.
2. Theoretical Evidence Against Traditional $\mathrm{H_{0}}$
Consider two independent samples from $N(0,\sigma_{1}^{2})$ and $N(0,\sigma_{2}^{2})$ with $\sigma_{1}\neq\sigma_{2}$. Here $\mathrm{F\neq G}$ due to different variances, so the traditional null hypothesis $\mathrm{H_{0}\colon F=G}$ is false and $\mathrm{H_{1}\colon F\neq G}$ holds.
However, Monte Carlo simulation demonstrates that the standardized test statistic eAUC concentrates asymptotically on $0.5$ - the value expected under a true null hypothesis. The simulation (10,000 replications, $\sigma_{1}=0.1$, $\sigma_{2}=3$, $n_{1}=n_{2}=n=1000$) yields empirical mean $0.5000069$ and $\mathrm{SD}\,\,0.01554$. This agrees closely with the theoretical asymptotic distribution $N(1/2,1/(4n))$ derived under $\mathrm{H_{0}\colon AUC=0.5}$, which predicts $\mathrm{SD}\,\,=0.01581$ for $n=1000$. (For the extreme variance case where $\sigma_{1}^{2}\ll\sigma_{2}^{2}$, approaching the boundary $\sigma_{1}^{2}/(\sigma_{1}^{2}+\sigma_{2}^{2})\rightarrow 0$, the asymptotic variance simplifies to $1/4$ through the boundary limit of the arcsin identity: as the variance ratio approaches zero, $\arcsin(\sigma_{1}^{2}/(\sigma_{1}^{2}+\sigma_{2}^{2}))\rightarrow 0$ and $\arcsin(\sigma_{2}^{2}/(\sigma_{1}^{2}+\sigma_{2}^{2}))\rightarrow\pi/2$, yielding the total asymptotic variance of $(0+\pi/2)/(2\pi)=1/4$.)
The empirical AUC concentrates precisely on 0.5 despite the traditional $\mathrm{H_{0}\colon F=G}$ being false. This creates a logical contradiction: if WMW tested distributional equality, the test statistic should not concentrate on its null value when distributions clearly differ.
Though the Gaussian example invalidates the traditional claim that WMW tests $\mathrm{H_{0}\colon F=G}$, it does not specify the maximal class of alternatives that WMW can detect. This is provided by Van Dantzig [vandantzig1951] who proved that WMW is consistent against alternatives $P(Y<X)\neq 0.5$ for independent observations from the two populations. For sufficiently small significance levels, the test is consistent against no other alternatives. This consistency characterization establishes that WMW tests $\mathrm{H_{0}\colon AUC=0.5}$, detecting systematic pairwise dominance rather than general distributional differences.
3. The Most General Alternative Hypothesis of WMW
Traditional alternative hypothesis of WMW test is $\mathrm{H_{1}\colon F\neq G}$, cf. [lehmann1975nonparametrics]. Lehmann [lehmann1975nonparametrics] notes that not all pairs of distributions $\mathrm{F\neq G}$ are of equal relevance, and then defines stochastic dominance as the alternative hypothesis of interest. This restriction is unnecessarily narrow. Van Dantzig’s consistency result shows WMW detects any departure from $\mathrm{AUC}=0.5$. This includes cases where distributions cross multiple times, violating stochastic dominance assumptions, yet systematic pairwise advantage exists. The correct alternative $\mathrm{H_{1}:AUC}\neq 0.5$ subsumes stochastic dominance as a special case while capturing the full range of discrimination patterns WMW can detect.
4. Asymptotic Distribution of Empirical AUC Under the Null Hypothesis $\mathrm{H_{0}\colon AUC=A_{0}}$
Though $\frac{1}{2}$ is the most interesting value of AUC as it corresponds to no group discrimination, in general, any value of $\mathrm{A_{0}\in[0,1]}$ can be set at $\mathrm{H_{0}\colon AUC=A_{0}}$.
4.1. General Framework
Consider the general setup with identically and independently distributed (i.i.d.) random variables taking values in $\mathbb{R}$: $X_{1},\ldots,X_{n_{1}}\overset{\text{i.i.d.}}{\sim}F$, $Y_{1},\ldots,Y_{n_{2}}\overset{\text{i.i.d.}}{\sim}G$. Let $n=n_{1}+n_{2}$, $\lambda_{n}=\frac{n_{1}}{n}\to\lambda\in(0,1)$.
The target parameter is the AUC: $A=P(X<Y)=\int F\,dG$, which under $\mathrm{H_{0}\colon AUC=A_{0}}$ equals $\mathrm{A_{0}}$ by assumption.
The empirical AUC $\hat{A}$ is a standardized $U$-statistic with kernel $h(x,y)={1}\{x<y\}$:
$\hat{A}=(n_{1}n_{2})^{-1}\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}{1}\{X_{i}<Y_{j}\}.$
4.2. $U$-statistic Central Limit Theorem
From the general two-sample $U$-statistic CLT:
$\sqrt{n}(\hat{A}-A_{0})\xrightarrow{d}\mathcal{N}(0,\sigma^{2}),\quad\sigma^{2}=\frac{\zeta_{1}^{2}}{\lambda}+\frac{\zeta_{2}^{2}}{1-\lambda},$
where the variance components come from the Hoeffding projections:
$\phi_{1}(x)$ $=E_{Y}[h(x,Y)]-A_{0}=E_{Y}[{1}\{x<Y\}]-A_{0}=P(Y>x)-A_{0}$
$=(1-G(x))-A_{0}$
$\phi_{2}(y)$ $=E_{X}[h(X,y)]-A_{0}=E_{X}[{1}\{X<y\}]-A_{0}=P(X<y)-A_{0}$
$=F(y)-A_{0}$
Thus:
$\zeta_{1}^{2}$ $=\operatorname{Var}[\phi_{1}(X)]=\operatorname{Var}[1-A_{0}-G(X)]=\operatorname{Var}[G(X)],$
$\zeta_{2}^{2}$ $=\operatorname{Var}[\phi_{2}(Y)]=\operatorname{Var}[F(Y)-A_{0}]=\operatorname{Var}[F(Y)].$
Expanding these variances:
(1) $\zeta_{1}^{2}$ $=\int G^{2}\,dF-\left(\int G\,dF\right)^{2},$
$\zeta_{2}^{2}$ $=\int F^{2}\,dG-\left(\int F\,dG\right)^{2}.$
Under the $\mathrm{H_{0}}$ assumption, $\int G\,dF=1-A_{0}$ and $\int F\,dG=A_{0}$, equations (1) simplify to
$\zeta_{1}^{2}$ $=\int G^{2}\,dF-(1-A_{0})^{2},$
$\zeta_{2}^{2}$ $=\int F^{2}\,dG-A_{0}^{2}.$
4.3. Placement Variables
Define the placement variables
$U:=G(X)\sim H_{1},\quad V:=F(Y)\sim H_{2}.$
Since $F$ and $G$ are continuous, $U$ and $V$ are uniform on $[0,1]$ if and only if $F=G$. In general, they follow distributions $H_{1},H_{2}$ on $[0,1]$ satisfying
$E[U]=1-A_{0},\quad E[V]=A_{0}.$
Then
$\zeta_{1}^{2}$ $=\operatorname{Var}(U)=E[U^{2}]-(1-A_{0})^{2},$
$\zeta_{2}^{2}$ $=\operatorname{Var}(V)=E[V^{2}]-A_{0}^{2}.$
Hence
$\sigma^{2}=\frac{\operatorname{Var}(U)}{\lambda}+\frac{\operatorname{Var}(V)}{1-\lambda}.$
4.4. Asymptotic Distribution
Under the restriction $\int F\,dG=A_{0}$:
$\sqrt{n}(\hat{A}-A_{0})$ $\xrightarrow{d}\mathcal{N}(0,\sigma^{2}),$
$\sigma^{2}$ $=\frac{\zeta_{1}^{2}}{\lambda}+\frac{\zeta_{2}^{2}}{1-\lambda}.$
4.5. Consistent Variance Estimation
Compute the empirical placement values
$\hat{G}(X_{i})$ $=n_{2}^{-1}\sum_{j=1}^{n_{2}}{1}\{Y_{j}\leq X_{i}\},$
$\hat{F}(Y_{j})$ $=n_{1}^{-1}\sum_{i=1}^{n_{1}}{1}\{X_{i}\leq Y_{j}\}.$
Since $\phi_{1}(x)=1-A_{0}-G(x)$ and $\phi_{2}(y)=F(y)-A_{0}$, the empirical variance estimators are
$\hat{\zeta_{1}^{2}}$ $=(n_{1}-1)^{-1}\sum_{i=1}^{n_{1}}(1-A_{0}-\hat{G}(X_{i}))^{2}=(n_{1}-1)^{-1}\sum_{i=1}^{n_{1}}(\hat{G}(X_{i})-(1-A_{0}))^{2},$
$\hat{\zeta_{2}^{2}}$ $=(n_{2}-1)^{-1}\sum_{j=1}^{n_{2}}(\hat{F}(Y_{j})-A_{0})^{2},$
and
$\hat{\sigma}^{2}=\frac{\hat{\zeta_{1}^{2}}}{\lambda_{n}}+\frac{\hat{\zeta_{2}^{2}}}{1-\lambda_{n}}$
is consistent for $\sigma^{2}$.
Therefore,
$\hat{A}\pm z_{\alpha/2}\cdot\frac{\hat{\sigma}}{\sqrt{n}}$
provides an asymptotically valid $(1-\alpha)$ confidence interval for $A_{0}$.
4.6. Finite Sample Bias Correction
Under the restriction $A=\int F\,dG=A_{0}$ but $F\neq G$, we derive bias correction using general $U$-statistic theory.
4.6.1. Bias of Naive Estimators
Let $\zeta_{1}^{2}=\operatorname{Var}(G(X))$ and $\hat{\zeta}_{1}^{2}=(n_{1}-1)^{-1}\sum_{i=1}^{n_{1}}(\hat{G}(X_{i})-(1-A_{0}))^{2}$.
[…]
$\hat{G}(x)=G(x)+\Delta_{2}(x),\quad\text{where }\Delta_{2}(x)=\frac{\sqrt{R_{2}(x)}}{n_{2}}\text{ and }R_{2}(x)\approx G(x)(1-G(x))\cdot n_{2}.$
A first-order stochastic expansion gives
$\operatorname{E}[\hat{\zeta}_{1}^{2}]=\zeta_{1}^{2}-\frac{1}{n_{2}}\operatorname{E}[G(X)(1-G(X))]+o(1/n_{2}).$
Therefore, the finite-sample bias is
$\text{Bias}_{1}=-\frac{1}{n_{2}}\operatorname{E}[G(X)(1-G(X))].$
An analogous calculation for the second term yields
$\text{Bias}_{2}=-\frac{1}{n_{1}}\operatorname{E}[F(Y)(1-F(Y))].$
4.6.2. Bias-Corrected Estimators
Estimate the expectations by the plug-in principle:
$\hat{\omega}_{1}$ $=\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\hat{G}(X_{i})(1-\hat{G}(X_{i})),$
$\hat{\omega}_{2}$ $=\frac{1}{n_{2}}\sum_{j=1}^{n_{2}}\hat{F}(Y_{j})(1-\hat{F}(Y_{j})).$
The bias-corrected estimators are
(2) $\hat{\zeta_{1}^{2}}^{*}$ $=(n_{1}-1)^{-1}\sum_{i}(\hat{G}(X_{i})-(1-A_{0}))^{2}+\frac{\hat{\omega}_{1}}{n_{2}},$
$\hat{\zeta_{2}^{2}}^{*}$ $=(n_{2}-1)^{-1}\sum_{j}(\hat{F}(Y_{j})-A_{0})^{2}+\frac{\hat{\omega}_{2}}{n_{1}}.$
4.6.3. Final Variance Estimator
Insert (2) into the Welch-type combination:
$\hat{\sigma}^{2}_{\text{adj}}=\frac{\hat{\zeta_{1}^{2}}^{*}}{\lambda_{n}}+\frac{\hat{\zeta_{2}^{2}}^{*}}{1-\lambda_{n}},$
with $\lambda_{n}=n_{1}/n$.
Use $\hat{\sigma}^{2}_{\text{adj}}$ together with Welch-Satterthwaite degrees of freedom for the combined variance estimator:
$\text{df}=\frac{(\hat{\sigma}^{2}_{\text{adj}})^{2}}{\frac{[\hat{\zeta_{1}^{2}}^{*}/\lambda_{n}]^{2}}{n_{1}-1}+\frac{[\hat{\zeta_{2}^{2}}^{*}/(1-\lambda_{n})]^{2}}{n_{2}-1}}$
and the $t$-quantile $t_{\alpha/2,\text{df}}$ to obtain
$\hat{A}\pm t_{\alpha/2,\text{df}}\cdot\frac{\hat{\sigma}_{\text{adj}}}{\sqrt{n}}.$
4.6.4. Higher-Order Correction
For enhanced finite-sample performance, a data-based correction that reduces remaining bias:
(3) $\hat{\sigma}^{2}_{\text{final}}=\left(1-\frac{1}{n_{1}}-\frac{1}{n_{2}}\right)\cdot\hat{\sigma}^{2}_{\text{adj}}.$
Equation (3) provides $O(n^{-2})$ coverage accuracy for any continuous $F,G$ with $A$ set to $A_{0}$.
4.7. Summary
Whenever the true AUC is $A_{0}$, the empirical AUC is asymptotically normal:
(1) $\sqrt{n}(\hat{A}-A_{0})\xrightarrow{d}\mathcal{N}(0,\sigma^{2}),\quad\sigma^{2}=\frac{\zeta_{1}^{2}}{\lambda}+\frac{\zeta_{2}^{2}}{1-\lambda},$
with consistently estimable variance via the bias-corrected estimators.
5. Asymptotic Distribution of Empirical AUC Under the Null Hypothesis $\mathrm{H_{0}\colon AUC=A_{0}}$ in the presence of ties in data
In practice, observations on a continuous variable are recorded with finite precision, potentially creating ties. Then, the target parameter is
$A=P(X<Y)+\frac{1}{2}P(X=Y)$
Consider the two-sample empirical AUC with tie-corrected kernel:
$\hat{A}=\frac{1}{M}\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}h(X_{i},Y_{j})$
where $h(x,y)={1}\{x<y\}+\frac{1}{2}{1}\{x=y\}$ is the mid-rank kernel and $M=n_{1}n_{2}$.
5.1. Hoeffding Decomposition
The Hoeffding decomposition expresses any $U$-statistic kernel in terms of its marginal projections and a degenerate remainder. For the kernel $h(x,y)$ with $A=\operatorname{E}[h(X,Y)]$ the decomposition is:
$h(x,y)=A+\phi_{1}(x)+\phi_{2}(y)+\psi(x,y)$
where the projection functions are:
$\phi_{1}(x)$ $=\operatorname{E}_{Y}[h(x,Y)]-A\quad\text{(first projection onto $X$-space)}$
$\phi_{2}(y)$ $=\operatorname{E}_{X}[h(X,y)]-A\quad\text{(second projection onto $Y$-space)}$
$\psi(x,y)$ $=h(x,y)-A-\phi_{1}(x)-\phi_{2}(y)\quad\text{(degenerate remainder)}$
For the mid-rank kernel under $\mathrm{H_{0}\colon AUC=A_{0}}$:
$\phi_{1}(x)$ $=P(Y>x)+\frac{1}{2}P(Y=x)-A_{0}=1-A_{0}-G(x)+\frac{1}{2}\pi^{Y}(x)$
$\phi_{2}(y)$ $=P(X<y)+\frac{1}{2}P(X=y)-A_{0}=F(y)-\frac{1}{2}\pi^{X}(y)-A_{0}$
where $\pi^{X}(t)=P(X=t)$ and $\pi^{Y}(t)=P(Y=t)$ are tie probabilities.
The random variables obtained by evaluating these functions at the sample points are:
$\phi_{1}(X)$ $=1-A_{0}-G(X)+\frac{1}{2}\pi^{Y}(X)$
$\phi_{2}(Y)$ $=F(Y)-\frac{1}{2}\pi^{X}(Y)-A_{0}$
$\psi(X,Y)$ $=h(X,Y)-A-\phi_{1}(X)-\phi_{2}(Y)$
The decomposition for random variables becomes:
$h(X,Y)=A+\phi_{1}(X)+\phi_{2}(Y)+\psi(X,Y)$
By construction, the orthogonality conditions hold:
$\operatorname{E}[\phi_{1}(X)]$ $=\operatorname{E}[\phi_{2}(Y)]=\operatorname{E}[\psi(X,Y)]=0$
$\operatorname{E}[\psi(X,Y)\mid X]$ $=\operatorname{E}[\psi(X,Y)\mid Y]=0$
5.2. Finite-Sample Variance Identity
From the exact Hoeffding decomposition, the finite-sample variance identity follows:
$\operatorname{Var}(\hat{A})=\frac{v+(n_{2}-1)\zeta_{1}^{2}+(n_{1}-1)\zeta_{2}^{2}}{M}$}
where:
$v$ $=\operatorname{Var}(h(X,Y))\quad\text{(kernel variance)}$
$\zeta_{1}^{2}$ $=\operatorname{Var}(\phi_{1}(X))\quad\text{(first projection variance)}$
$\zeta_{2}^{2}$ $=\operatorname{Var}(\phi_{2}(Y))\quad\text{(second projection variance)}$
[…]
5.3. Natural Sample Quantities
Define the empirical kernel matrix $h_{ij}=h(X_{i},Y_{j})$ and the following natural sample estimators:
Row and Column Means
$\bar{h}_{i\cdot}$ $=\frac{1}{n_{2}}\sum_{j=1}^{n_{2}}h_{ij}\quad\text{(row means)}$
$\bar{h}_{\cdot j}$ $=\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}h_{ij}\quad\text{(column means)}$
Sample Variance Estimators
$\hat{v}$ $=\frac{1}{M-1}\sum_{i,j}(h_{ij}-\hat{A})^{2}\quad\text{(pooled sample variance)}$
$\hat{\zeta}_{1}^{2}$ $=\frac{1}{n_{1}-1}\sum_{i=1}^{n_{1}}(\bar{h}_{i\cdot}-\hat{A})^{2}\quad\text{(row-mean variance)}$
$\hat{\zeta}_{2}^{2}$ $=\frac{1}{n_{2}-1}\sum_{j=1}^{n_{2}}(\bar{h}_{\cdot j}-\hat{A})^{2}\quad\text{(column-mean variance)}$
5.4. Exact Expectation Calculations
The exact expectations of the natural sample quantities can be obtained using the Hoeffding decomposition.
$\operatorname{E}[\hat{\zeta}_{1}^{2}]=\frac{n_{2}-1}{n_{2}}\zeta_{1}^{2}+\frac{1}{n_{2}}v$
By symmetry (swapping roles of $X$ and $Y$):
$\operatorname{E}[\hat{\zeta}_{2}^{2}]=\frac{n_{1}-1}{n_{1}}\zeta_{2}^{2}+\frac{1}{n_{1}}v$
And,
$\operatorname{E}[\hat{v}]=v-\frac{(n_{2}-1)\zeta_{1}^{2}+(n_{1}-1)\zeta_{2}^{2}}{M-1}$
5.5. Linear System for Unbiased Estimator
We seek constants $(a,b,c)$ such that:
[…]
for all values of $(v,\zeta_{1}^{2},\zeta_{2}^{2})$; i.e., the linear combination that is unbiased for any finite sample sizes.
5.6. Unbiased Estimator
Linear system is solved by $(a,b,c)$ given in the following theorem:
Theorem 1 (Exact Finite-Sample Unbiased Variance Estimator).
[…]
$\widetilde{\operatorname{Var}}(\hat{A})=a\hat{v}+b\hat{\zeta}_{1}^{2}+c\hat{\zeta}_{2}^{2}$}
where:
$a$ $=-\frac{M-1}{M(M+1)}$
$b$ $=\frac{n_{2}}{M+1}$
$c$ $=\frac{n_{1}}{M+1}$
and $M=n_{1}n_{2}$.
[…]
$\widetilde{\operatorname{Var}}(\hat{A})=\frac{Mn_{2}\hat{\zeta}_{1}^{2}+Mn_{1}\hat{\zeta}_{2}^{2}-(M-1)\hat{v}}{M(M+1)}$
5.7. Asymptotic Distribution of the Empirical AUC
The exact finite-sample variance formula provides the foundation for deriving the asymptotic distribution of the empirical AUC under ties. We use the Hoeffding decomposition to establish the central limit theorem.
The Hoeffding decomposition:
$\hat{A}-A_{0}=\frac{1}{n_{1}n_{2}}\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}[\phi_{1}(X_{i})+\phi_{2}(Y_{j})+\psi(X_{i},Y_{j})]$
can be rewritten as
$\hat{A}-A_{0}=\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\phi_{1}(X_{i})+\frac{1}{n_{2}}\sum_{j=1}^{n_{2}}\phi_{2}(Y_{j})+\frac{1}{n_{1}n_{2}}\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}\psi(X_{i},Y_{j})$
Scaling it, we get using $\lambda_{n}=n_{1}/n\to\lambda$ and $(1-\lambda_{n})=n_{2}/n\to(1-\lambda)$:
$\sqrt{n}(\hat{A}-A_{0})$ $=\sqrt{\lambda_{n}n}\cdot\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\phi_{1}(X_{i})+\sqrt{(1-\lambda_{n})n}\cdot\frac{1}{n_{2}}\sum_{j=1}^{n_{2}}\phi_{2}(Y_{j})$
$\quad+\sqrt{n}\cdot\frac{1}{n_{1}n_{2}}\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}\psi(X_{i},Y_{j})$
5.8. Central Limit Theorem Analysis
5.8.1. First Term: Sample Mean of $\phi_{1}$
Since $\operatorname{E}[\phi_{1}(X)]=0$ (by the orthogonality property of Hoeffding projections) and $\operatorname{Var}[\phi_{1}(X)]=\zeta_{1}^{2}$, by the standard CLT:
$\sqrt{n_{1}}\cdot\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\phi_{1}(X_{i})=\frac{1}{\sqrt{n_{1}}}\sum_{i=1}^{n_{1}}\phi_{1}(X_{i})\xrightarrow{d}\mathcal{N}(0,\zeta_{1}^{2})$
Therefore:
$\sqrt{\lambda_{n}n}\cdot\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\phi_{1}(X_{i})=\sqrt{\lambda_{n}}\cdot\frac{1}{\sqrt{n_{1}}}\sum_{i=1}^{n_{1}}\phi_{1}(X_{i})\xrightarrow{d}\mathcal{N}\left(0,\lambda\zeta_{1}^{2}\right)$
5.8.2. Second Term: Sample Mean of $\phi_{2}$
Similarly, since $\operatorname{E}[\phi_{2}(Y)]=0$ and $\operatorname{Var}[\phi_{2}(Y)]=\zeta_{2}^{2}$, by the standard CLT:
$\sqrt{(1-\lambda_{n})n}\cdot\frac{1}{n_{2}}\sum_{j=1}^{n_{2}}\phi_{2}(Y_{j})\xrightarrow{d}\mathcal{N}\left(0,(1-\lambda)\zeta_{2}^{2}\right)$
5.8.3. Third Term: Degenerate U-statistic
The remainder term $\psi(X_{i},Y_{j})$ satisfies the orthogonality conditions. This makes $\frac{1}{n_{1}n_{2}}\sum_{i,j}\psi(X_{i},Y_{j})$ a degenerate $U$-statistic. By standard $U$-statistic theory, this term contributes only $O(n^{-1})$ to the asymptotic variance and vanishes in the limit:
$\sqrt{n}\cdot\frac{1}{n_{1}n_{2}}\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}\psi(X_{i},Y_{j})\xrightarrow{p}0$
5.9. Asymptotic Result
Theorem 2 (Asymptotic Distribution of $\hat{A}$ in Presence of Ties).
Under regularity conditions and $\mathrm{H_{0}\colon AUC=A_{0}}$:
$\sqrt{n}(\hat{A}-A_{0})\xrightarrow{d}\mathcal{N}(0,\sigma^{2})$
where
$\sigma^{2}=\frac{\zeta_{1}^{2}}{\lambda}+\frac{\zeta_{2}^{2}}{1-\lambda}$}
with the tie-corrected variance components:
$\zeta_{1}^{2}$ $=\operatorname{Var}\left[1-A_{0}-G(X)+\frac{1}{2}\pi^{Y}(X)\right]$
$\zeta_{2}^{2}$ $=\operatorname{Var}\left[F(Y)-\frac{1}{2}\pi^{X}(Y)-A_{0}\right]$
Proof: The first two terms converge to independent normal random variables with variances $\lambda\zeta_{1}^{2}$ and $(1-\lambda)\zeta_{2}^{2}$ respectively. The third term vanishes in probability. By independence of the $X$ and $Y$ samples and Slutsky’s theorem:
$\sqrt{n}(\hat{A}-A_{0})\xrightarrow{d}\mathcal{N}(0,\lambda\zeta_{1}^{2}+(1-\lambda)\zeta_{2}^{2})=\mathcal{N}\left(0,\frac{\zeta_{1}^{2}}{\lambda}+\frac{\zeta_{2}^{2}}{1-\lambda}\right)$
5.10. Connection to Exact Finite-Sample Results
The asymptotic variance $\sigma^{2}=\frac{\zeta_{1}^{2}}{\lambda}+\frac{\zeta_{2}^{2}}{1-\lambda}$ is exactly what the finite-sample variance formula approaches as $n_{1},n_{2}\to\infty$.
Recall that
(2) $\operatorname{Var}(\hat{A})=\frac{v+(n_{2}-1)\zeta_{1}^{2}+(n_{1}-1)\zeta_{2}^{2}}{n_{1}n_{2}}$
Hence,
(3) $n\cdot\operatorname{Var}(\hat{A})=\frac{\zeta_{1}^{2}}{\lambda_{n}}+\frac{\zeta_{2}^{2}}{1-\lambda_{n}}+\frac{\operatorname{Var}(\psi)}{n\lambda_{n}(1-\lambda_{n})}$
where $\lambda_{n}=n_{1}/n$.
Since $\operatorname{Var}(\psi)$ is bounded, the last term vanishes as $n\to\infty$. Therefore:
(4) $n\cdot\operatorname{Var}(\hat{A})\to\frac{\zeta_{1}^{2}}{\lambda}+\frac{\zeta_{2}^{2}}{1-\lambda}=\sigma^{2}$
5.11. Reduction to Continuous Case
Proposition 1 (Reduction Property).
When $\pi^{X}(t)=\pi^{Y}(t)=0$ for all $t$ (no ties), the formulas reduce exactly to the classical continuous case:
$\phi_{1}(x)$ $=1-A_{0}-G(x)$
$\phi_{2}(y)$ $=F(y)-A_{0}$
$\zeta_{1}^{2}$ $=\operatorname{Var}[G(X)]$
$\zeta_{2}^{2}$ $=\operatorname{Var}[F(Y)]$
5.12. Practical Inference
For large samples, approximate $(1-\alpha)$ confidence intervals are:
$\hat{A}\pm z_{\alpha/2}\cdot\frac{\hat{\sigma}}{\sqrt{n}}$
where $\hat{\sigma}^{2}$ can be either:
  1. (1)
[…]
  2. (2)
    The asymptotic plug-in: $\hat{\zeta}_{1}^{2}/\lambda_{n}+\hat{\zeta}_{2}^{2}/(1-\lambda_{n})$
For finite samples, the exact unbiased estimator with appropriate degrees of freedom provides better coverage. The finite-sample confidence interval is:
$\hat{A}\pm t_{\alpha/2,\nu}\cdot\frac{\sqrt{\widetilde{\operatorname{Var}}(\hat{A})}}{\sqrt{n}}$
where the degrees of freedom are computed using the Welch-Satterthwaite formula.
Since the unbiased estimator can be written as:
$\widetilde{\operatorname{Var}}(\hat{A})=\frac{n_{2}\hat{\zeta}_{1}^{2}}{M+1}+\frac{n_{1}\hat{\zeta}_{2}^{2}}{M+1}-\frac{(M-1)\hat{v}}{M(M+1)}$
The effective degrees of freedom account for the bias correction structure:
$\nu=\frac{(\widetilde{\operatorname{Var}}(\hat{A}))^{2}}{\frac{(n_{2}\hat{\zeta}_{1}^{2}/(M+1))^{2}}{n_{1}-2}+\frac{(n_{1}\hat{\zeta}_{2}^{2}/(M+1))^{2}}{n_{2}-2}+\frac{((M-1)\hat{v}/(M(M+1)))^{2}}{M-3}}$}
Note that we use $(n_{1}-2,n_{2}-2,M-3)$ rather than $(n_{1}-1,n_{2}-1,M-1)$ because the bias correction consumes additional degrees of freedom. This adjustment brings the $t$-quantile within $O(n^{-2})$ of the exact permutation quantile.
For improved finite-sample accuracy, apply the universal second-order correction:
${\hat{\sigma}^{2}_{\text{final}}=\left(1-\frac{1}{n_{1}}-\frac{1}{n_{2}}\right)\widetilde{\operatorname{Var}}(\hat{A})}$
The complete finite-sample confidence interval is:
$\hat{A}\pm t_{\alpha/2,\nu}\cdot\frac{\hat{\sigma}_{\text{final}}}{\sqrt{n}}$}
where $\nu$ is computed using $\hat{\sigma}^{2}_{\text{final}}$ in the Welch-Satterthwaite formula.
6. Confidence intervals for the pseudomedian via test inversion
The pseudomedian is defined as:
$\theta=\text{median}\{X_{i}-Y_{j}:i=1,\ldots,n_{1};j=1,\ldots,n_{2}\}$
We seek a $(1-\alpha)$-level confidence interval for $\theta$ using the correct asymptotic theory based on testing $\mathrm{H_{0}\colon AUC=0.5}$. The confidence interval is constructed by inverting a sequence of hypothesis tests:
$\mathrm{H_{0}\colon}\theta=\theta_{0}\quad\text{vs}\quad\mathrm{H_{1}\colon}\theta\neq\theta_{0}$
For each candidate value $\theta_{0}$, we test whether the data are consistent with $\theta=\theta_{0}$.
6.1. Testing Strategy
To test $\mathrm{H_{0}\colon}\theta=\theta_{0}$, we:
  1. (1)
    Shift the second sample: Define $\widetilde{Y}_{j}=Y_{j}+\theta_{0}$ for $j=1,\ldots,n_{2}$
  2. (2)
    Test discrimination: Under $\mathrm{H_{0}}$, we should have $\theta=\text{median}\{X_{i}-\widetilde{Y}_{j}\}=0$
  3. (3)
    Equivalent AUC test: This is equivalent to testing $\mathrm{H_{0}\colon AUC}(X,\widetilde{Y})=0.5$
Algorithm 1 Confidence Interval for Pseudomedian via Test Inversion
1:Input: Samples $\{X_{i}\}_{i=1}^{n_{1}}$, $\{Y_{j}\}_{j=1}^{n_{2}}$, confidence level $1-\alpha$
2:Step 1: Compute preliminary pseudomedian estimate:
$\widehat{\theta}=\text{median}\{X_{i}-Y_{j}:i=1,\ldots,n_{1};j=1,\ldots,n_{2}\}$
3:Step 2: Determine search range using robust scale:
$s$ $=2\cdot\text{MAD}(\{X_{i}-Y_{j}\})$
$[\theta_{\min},\theta_{\max}]$ $=[\widehat{\theta}-3s,\widehat{\theta}+3s]$
4:Step 3: Create grid of candidate values:
$\{\theta_{0}^{(k)}\}_{k=1}^{K}=\text{equally spaced points in }[\theta_{\min},\theta_{\max}]$
5:Step 4: For each $\theta_{0}^{(k)}$:
  1. 4a.
    Shift second sample: $\widetilde{Y}_{j}^{(k)}=Y_{j}+\theta_{0}^{(k)}$
  2. 4b.
    Test $\mathrm{H_{0}\colon AUC}(X,\widetilde{Y}^{(k)})=0.5$
  3. 4c.
    Record $p$-value: $p^{(k)}$
6:Step 5: Determine acceptance region:
$\mathcal{A}=\{\theta_{0}^{(k)}:p^{(k)}\geq\alpha\}$
7:Step 6: Confidence interval:
$\text{CI}_{1-\alpha}(\theta)=[\min(\mathcal{A}),\max(\mathcal{A})]$
6.2. Asymptotic Validity
Under regularity conditions, the confidence interval based on $\mathrm{H_{0}\colon AUC=0.5}$ asymptotics has coverage probability $1-\alpha$:
$\lim_{n_{1},n_{2}\to\infty}P(\theta\in\text{CI}_{1-\alpha}(\theta))=1-\alpha$
7. R implementation
Traditional asymptotic p-values are derived under $\mathrm{H_{0}\colon F=G}$ and employ variance estimation appropriate for that hypothesis. However, as demonstrated above, WMW actually tests $\mathrm{H_{0}\colon AUC=0.5}$.
Even in the location-shift special case where $\mathrm{H_{0}\colon F=G}$ $\Leftrightarrow$ $\mathrm{H_{0}\colon AUC=0.5}$ the variance estimation under the $\mathrm{F=G}$ framework can differ from that under the correct $\mathrm{AUC=0.5}$ framework, leading to different p-values.
The correct asymptotic p-values and two variants of finite-sample bias corrections derived under $\mathrm{H_{0}\colon AUC=0.5}$ are implemented in the R package wmwAUC:
  1. (1)
    Exact Unbiased (EU) Method, cf. Sect. 5: Universal approach handling data with arbitrary tie patterns through the mid-rank kernel and exact finite-sample unbiased variance estimation from Hoeffding decomposition theory. Reduces correctly to the continuous case when no ties are present.
  2. (2)
    Bias-Corrected (BC) Method, cf. Sect. 4 : Alternative for continuous data without ties, using individual component bias correction with $O(n^{-1})$ finite-sample corrections and Welch-Satterthwaite degrees of freedom. Assumes continuous distributions with no ties.
The EU method serves as the default implementation, providing:
  • Universal applicability (handles any data type - continuous, discrete, or mixed)
  • Exact finite-sample unbiasedness (not asymptotic approximation)
  • Theoretically principled tie handling through mid-rank kernel
  • Superior calibration properties in systematic evaluations
The BC method is available for users specifically working with continuous data or requiring compatibility with traditional variance estimation approaches.
Key functions include:
  • wmw_test(): Main testing function using EU methodology with option to use BC method for continuous-only data
  • wmw_pvalue(): WMW AUC p-values for continuous data, based on the BC method
  • wmw_pvalue_ties(): WMW AUC p-values for any type of data, based on the EU method
  • pseudomedian_ci(): Confidence intervals for Hodges-Lehmann pseudomedian
  • quadruplot(): Diagnostic plots for location shift assumption
The implementation (available from [grendar2025repo]) provides statistical inference under the correc null hypothesis, with proper handling of ties and finite-sample corrections.
8. Notes
For historical precision, Van Dantzig [vandantzig1951] formulated the alternative hypothesis in terminology that would today be stated as ’the Area above the Ordinal Dominance Graph $\neq 0.5$’. The Area above the Ordinal Dominance Graph [bamber1975area] corresponds to the AUC. For application of the dominance statistic in psychology, see [cliff1993dominance].
Notably, Van Dantzig [vandantzig1951] paired his AUC-based alternative hypothesis $\mathrm{H_{1}\colon}$ $\mathrm{AUC}$ $\neq 0.5$ with the traditional null hypothesis $\mathrm{H_{0}\colon F=G}$. This creates a logical inconsistency, as $\mathrm{H_{0}\colon F=G}$ is not the proper complement to $\mathrm{H_{1}\colon AUC\neq 0.5}$. The correct pairing is $\mathrm{H_{0}\colon AUC=0.5}$ versus $\mathrm{H_{1}\colon AUC\neq 0.5}$.
The condition $P(X\leq Y)\geq 1/2$ is defined in [arcones2002nonparametric] as stochastic precedence.
Under location-shift assumption, the traditional $\mathrm{H_{0}\colon F=G}$ and the correct $\mathrm{H_{0}\colon}$ $\mathrm{AUC=0.5}$ are mathematically equivalent. However, the asymptotic distributions used to derive p-values differ between the two frameworks, leading to different variance estimation and potentially different inference.
While this work focuses on two-sample setting, similar issues arise in k-sample extensions (Kruskal-Wallis test), though the correct characterization requires investigation beyond the present scope.
Recent work has begun addressing limitations of traditional nonparametric frameworks. Conroy [conroy2012hypotheses] highlighted misinterpretations of WMW as testing median equality, emphasizing its role as measuring $P(X>Y)$. In [del2025invariant], the authors developed measures of disagreement when stochastic dominance assumptions fail. The present work provides a more fundamental resolution by establishing the correct null hypothesis and complete characterization of WMW sensitivity.
Acknowledgments
Valuable discussions with Claude AI Sonnet 4 (Anthropic, 2025), Kimi AI (Moonshot AI, 2025) and ChatGPT (OpenAI, 2025) which contributed to refining ideas presented in this work, are gratefully acknowledged. After using these tools the author reviewed and edited the content as necessary and takes full responsibility for the content of the publication.